The table below shows the monthly demand schedule for a good in a duopoly market. The two producers in this market each faces $5,000 of fixed costs per month. There are no marginal costs.

Quantity Price ($) TR ($) MR ($)
0 40 0
200 35 7,000 35
400 30 12,000 25
600 25 15,000 15
800 20 16,000 5
1,000 15 15,000 −5
1,200 10 12,000 −15
1,400 5 7,000 −25
1,600 0 0 −35


Instructions: Enter your answers to the nearest whole number.

a. If they evenly split the quantity a monopolist would produce, the monthly profit for each duopolist is $__.

.

b. Suppose duopolist A decides to increase production by 200 units.

Duopolist A will now produce __ units and charge a price of $__.

Duopolist B will now produce __ units and charge a price of $ __.

  The monthly profit for duopolist A is $ __ and for duopolist B $__.
Answer #1

a) Setting MR=MC

A monopolist would produce where the TR is maximized so each duopolist will produce 800/2 =400 units at a price of 20

Profit = TR-TC

400*20 - 5000 = 3000

b) When Duopolist A increases production by 200 units

Duopolist A will now produce 400+200 = 600 units and charge a price of $15

Duopolist B will now produce 400 units and charge a price of $15

The monthly profit for duopolist A is 600*15-5000 = 4000  and for duopolist, B is 400*15-5000 = 1000